### v2

# Quantum dispersionless KdV hierarchy revisited

Abstract.
We quantize Hamiltonian structures with hydrodynamic leading terms using the Heisenberg vertex algebra. As an application, we construct the quantum dispersionless KdV hierarchy via a non-associative Weyl quantization and compute the corresponding eigenvalue problem.
1. Introduction
Since the proof of the Witten-Kontsevich theorem [15, 23], integrable hierarchies have provided important tools for studying various problems in mathematical physics. For example, in 2d topological field theories, corresponding partition functions are controlled by evolutionary PDEs [1, 6]. In symplectic field theories [7], partition functions satisfy a family of Schrödinger type equations, and the commuting quantum Hamiltonian operators can be viewed as quantum integrable hierarchies. While the relationship between 2d topological field theories and evolutionary integrable PDEs have been studied in great detail during last three decades, the interplay between symplectic field theories and quantum integrable hierarchies are still largely unknown.
In [2], a quantization framework is established for arbitrary cohomological field theories, where quantum Hamiltonian structures are given by the canonical quantization of the original Poisson structures of hydrodynamic type, and quantum Hamiltonians are given by intersection numbers and Hodge integrals over double ramification cycles. Let us take the trivial one-dimensional cohomological field theory as an example (see [2, 5, 7]). The genus zero (classical) integrable hierarchy is the dispersionless KdV hierarchy which admits the following Hamiltonian form:
$\frac{\partial v}{\partial t_{n}}=\frac{v^{n}}{n!}\frac{\partial v}{\partial x}=\frac{\partial}{\partial x}\left(\frac{\delta H_{n+1}^{c}}{\delta v}\right),\quad H_{n}^{c}=\int\frac{v^{n+1}}{(n+1)!},\quad n\geq 0,$
and the corresponding Hamiltonian structure is given by the Poisson bracket of the form
(1.1) $\{v(x),v(y)\}=\delta^{\prime}(x-y).$
Note that the Hamiltonians are in involution:
$\{H_{n}^{c},H_{m}^{c}\}=0,\quad n,m\geq 0.$
To quantize this Poisson bracket, we expand the field $v(x)$ into Fourier modes
$v(x)=\sum_{k\in\mathbb{Z}}v_{k}e^{ikx},$
then the Poisson bracket above is equivalent to
$\{v_{m},v_{n}\}=i\,m\,\delta_{m+n,0}.$
By canonical quantization, for $n\geq 1$, let us denote by $\hat{q}_{n}$ the creators corresponding to modes $v_{n}$ and $\hat{p}_{n}$ the annihilators corresponding to modes $v_{-n}$. Then the Poisson bracket is quantized into the commutators
(1.2) $[\hat{p}_{n},\hat{q}_{m}]=\hbar\,n\,\delta_{m,n},\quad[\hat{p}_{n},\hat{p}_{m}]=[\hat{q}_{n},\hat{q}_{m}]=0,\quad n,m\geq 1,$
where $\hbar$ is a formal parameter, and the Hamiltonians $H_{n}^{c}$ is quantized into
$\hat{H}_{n}^{c}=\int:\frac{v^{n+1}}{(n+1)!}:.$
To explicitly compute these quantum Hamiltonians, we first substitute
$v=\sum_{n\geq 1}\hat{q}_{n}e^{inx}+\hat{p}_{n}e^{-inx},$
and normally order these operators by putting annihilators $\hat{p}_{n}$ on the right to obtain
$:\frac{v^{n+1}}{(n+1)!}:\,=\sum_{k\in\mathbb{Z}}f_{k}^{(n)}(\hat{q}_{i},\hat{p}_{j})e^{ikx},$
then we set $\hat{H}_{n}^{c}=f_{0}^{(n)}$. For example, we see that
$\hat{H}_{1}^{c}=\sum_{n\geq 1}\hat{q}_{n}\hat{p}_{n},\quad\hat{H}_{2}^{c}=\frac{1}{2}\sum_{m,n\geq 1}\hat{q}_{n}\hat{q}_{m}\hat{p}_{m+n}+\hat{q}_{m+n}\hat{p}_{n}\hat{p}_{m}.$
One can verify that classically commutative Hamiltonians $H_{n}^{c}$ do not quantize into commutative operators $\hat{H}_{n}^{c}$ under the commutators defined in (1.2). It is observed by Eliashberg that there exist quantum corrections to these operators such that they become mutually commuting operators after corrections. These corrected quantum Hamiltonians are given by the generating series (see [5, 20])
$\hat{H}^{q}(z)=\frac{1}{s(\sqrt{\hbar}z)}\int:\exp\left(zs(i\sqrt{\hbar}z\partial_{x})v(x)\right):,\quad\hat{H}^{q}(z)=1+\sum_{n\geq 0}\hat{H}_{n}^{q}z^{n+1},$
where the function
$s(t)=\frac{e^{t/2}-e^{-t/2}}{t}=\sum_{n\geq 0}\frac{t^{2n}}{4^{n}(2n+1)!}.$
Then it is proved that $\hat{H}_{n}^{q}$ is a correction of $\hat{H}_{n}^{c}$ in the sense that $\hat{H}_{n}^{q}|_{\hbar=0}=\hat{H}_{n}^{c}$ and
$\left[\hat{H}_{n}^{q},\hat{H}_{m}^{q}\right]=0,\quad n,m\geq 1.$
In [2], it is proved that these quantum corrections can be obtained by computing Hodge integrals over double ramification cycles. Furthermore, the corresponding symplectic field theoretic potential is computed in [5]. More aspects of the spectral problem are studied in [10, 11, 21].
In this paper, we approach the quantization problem of the dispersionless KdV hierarchy in the view point of vertex algebra. We obtain a slightly different deformation quantization of the Hamiltonian structure (1.1) from that obtained in [2], and we find a natural combinatorial expression for the corrected quantum Hamiltonians, which indicates a deep relationship between quantum integrable systems and vertex algebra.
To present the main results of the paper, we first introduce some notations. Let us denote by
$\mathcal{A}=\mathbb{C}\left[v,v^{(1)},v^{(2)},\dots\right]$
the ring of differential polynomial and by $\partial_{x}$ the derivation on $\mathcal{A}$ defined by
$\partial_{x}=\sum_{n\geq 0}v^{(n+1)}\frac{\partial}{\partial v^{(n)}},\quad v^{(0)}:=v.$
Using this notation, we can formally view $v$ as a function $v(x)$ with the independent variable $x$, then we have $v^{(n)}=\partial_{x}^{n}v(x)$. Denote by $\mathcal{F}=\mathcal{A}/\partial_{x}\mathcal{A}$ the space of local functionals, and for an element $f\in\mathcal{A}$, denote by $\int f$ its class in $\mathcal{F}$. These spaces play very important roles in the theory of integrable hierarchies. For example, the Hamiltonian structure (1.1) can be written as a Lie bracket on $\mathcal{F}$:
(1.3) $\left\{\int f,\int g\right\}=\int\frac{\delta f}{\delta v}\partial_{x}\frac{\delta g}{\delta v},\quad\frac{\delta f}{\delta v}:=\sum_{n\geq 0}(-\partial_{x})^{n}\frac{\partial f}{\partial v^{(n)}}.$
Denote by $V$ the Heisenberg vertex algebra generated by the field $b(z)$, whose modes form the Heisenberg Lie algebra:
$b(z)=\sum_{n\in\mathbb{Z}}b_{(n)}z^{-n-1},\quad\left[b_{(n)},b_{(m)}\right]=\hbar\,n\,\delta_{m+n,0}.$
Let $|0\rangle\in V$ be the vacuum state, then the state space $V$ is a free $\mathbb{C}[\hbar]$-module spanned by states of the form
$b_{(-k_{1}-1)}\dots b_{(-k_{n}-1)}|0\rangle,\quad k_{i}\geq 0,\quad n\geq 0.$
Hence, we have a natural $\hbar$-linear isomorphism of vector space
$\varphi:\mathcal{A}[\hbar]\to V,\quad v^{(k_{1})}\dots v^{(k_{n})}\mapsto k_{1}!\dots k_{n}!\,b_{(-k_{1}-1)}\dots b_{(-k_{n}-1)}|0\rangle.$
[…]
Theorem 1.1.
The Lie bracket
(1.4) $[F,G]=\varphi^{-1}\left(\varphi(G)_{(0)}\varphi(F)\right),\quad F,G\in\mathcal{F}[\hbar],$
defines a deformation quantization of the Hamiltonian structure (1.3). In terms of differential polynomials, this bracket is of the form
$[F,G]=\sum_{m\geq 1}\frac{\hbar^{m}}{m!}\sum_{\begin{subarray}{c}r_{1},\dots,r_{m}\geq 0\\ s_{1},\dots,s_{m}\geq 0\end{subarray}}$ $\frac{\partial^{m}f}{\partial v^{(s_{1})}\dots\partial v^{(s_{m})}}\frac{(-1)^{\sum r_{i}}\prod(r_{i}+s_{i}+1)!}{(2m-1+\sum r_{i}+\sum s_{i})!}$
$\cdot\partial_{x}^{2m-1+\sum r_{i}+\sum s_{i}}\frac{\partial^{m}g}{\partial v^{(r_{1})}\dots\partial v^{(r_{m})}},\quad F,G\in\mathcal{F},$
here $f,g\in\mathcal{A}$ are arbitrary lifts of $F$ and $G$, respectively.
Next, let us construct the quantization of the Hamiltonians of dispersionless KdV hierarchy. By abusing the notation, we use $b$ to denote the state $b_{(-1)}|0\rangle$. We then define the quantum Hamiltonian densities $h_{n}\in\mathcal{A}[\hbar]$ by the recursion relation
$h_{0}=\varphi^{-1}(b),\quad h_{n}=\frac{1}{n}\sum_{k=0}^{n-1}\frac{1}{\binom{n+1}{k+1}}\varphi^{-1}\left(:\varphi(h_{k})\varphi(h_{n-k-1}):\right),\quad n\geq 1.$
We call this a non-associative Weyl quantization since each $h_{n}$ corresponds to the state obtained by averaging all the possible ways of associating $n+1$ states of $b$. See Sect. 3.1 for details and a combinatorial way for constructing $h_{n}$. Though these Hamiltonian densities are defined via vertex algebra, they satisfy a straightforward recursion relation as differential polynomials.
Proposition 1.2.
We have $h_{0}=v$ and for $n\geq 1$
(1.5) $h_{n}=\frac{1}{n(n+1)}$ $\left(\sum_{s,t\geq 0}\frac{(s+t+1)!}{s!t!}v^{(s)}v^{(t)}\frac{\partial}{\partial v^{(s+t)}}\right.$
$\left.+\hbar\sum_{s,t\geq 0}\frac{(s+1)!(t+1)!}{(s+t+2)!}v^{(s+t+2)}\frac{\partial^{2}}{\partial v^{(s)}\partial v^{(t)}}\right)h_{n-1}.$
From this recursion we see that $h_{n}$ is indeed a deformation of the classical Hamiltonians in the sense that
$\lim_{\hbar\to 0}h_{n}=\frac{v^{n+1}}{(n+1)!}.$
Theorem 1.3.
Denote by $H_{n}=\int h_{n}$, then
$\left[H_{n},H_{m}\right]=0,\quad n,m\geq 0,$
here the bracket is the deformed quantum Hamiltonian structure defined in (1.4). Hence, the quantum Hamiltonians $H_{n}$ form a quantization of the dispersionless KdV hierarchy.
Finally, let us consider the eigenvalue problem. Due to our definition of $h_{n}$, it is natural to consider the operators $\varphi(h_{n})_{(n)}\in\mathrm{End}(V)$.
Remark 1.1.
Note that $\varphi(h_{n})$ is of conformal weight $n+1$, hence the operator $\varphi(h_{n})_{(n)}$ is of conformal weight 0.
Proposition 1.4.
Let $f\in\mathcal{A}$ be a differential polynomial. Then for $n\geq 0$, we have
[…]
$\cdot\partial_{x}^{2m-1-n+\sum r_{i}+\sum s_{i}}\left(\frac{\partial^{m}f}{\partial v^{(r_{1})}\dots\partial v^{(r_{m})}}\right)\frac{\partial^{m}}{\partial v^{(s_{1})}\dots\partial v^{(s_{m})}}$
[…]
$S_{\lambda}=\det{\left({\bf{e}}_{\lambda_{i}-i+j}\right)_{1\leq i,j\leq\ell(\lambda)}},$
here $\lambda=(\lambda_{1},\dots,\lambda_{\ell(\lambda)})$ is a partition of length $\ell(\lambda)$, and ${\bf{e}}_{n}\in\mathcal{A}[\sqrt{\hbar}]$ is given by
$\sum_{n\in\mathbb{Z}}{\bf{e}}_{n}z^{n}=\exp\left(\sum_{k\geq 0}\frac{\hbar^{k/2}v^{(k)}}{(k+1)!}z^{k+1}\right).$
Remark 1.2.
One needs to introduce $\sqrt{\hbar}$ to consider the eigenvalue problem.
Theorem 1.5.
Schur polynomials serve as a complete set of common eigenvectors of operators $\hat{H}_{n}$. More precisely, we have
$\hat{H}_{n}S_{\lambda}=\frac{\hbar^{\frac{n+1}{2}}}{n+1}\sum_{k=0}^{n}\binom{n+1}{k}\left(\binom{-\ell(\lambda)}{k+1}+\sum_{i=1}^{\ell(\lambda)}\binom{\lambda_{i}-i}{k}\right)S_{\lambda}.$
As a byproduct, we also prove an explicit formula of $h_{n}$ represented by Schur polynomials.
Proposition 1.6.
We have
$h_{n}=\frac{1}{n+1}\sum_{k=0}^{n}\frac{1}{\binom{n}{k}}S_{(n-k+1,\underbrace{1,\dots,1}_{k})}.$
Organization of the paper
This paper is organized as follows. In Sect. 2, after recalling some basic facts about vertex algebras, we prove Theorem 1.1 and Proposition 1.4 in a slightly more general setting. In Sect. 3 we construct a quantization of the dispersionless KdV hierarchy and prove various properties of the quantum Hamiltonians. Our main tool for the proofs is based on the work [18] studying actions of $W$-type operators on Schur functions. Finally, in Sect. 4, we give some concluding remarks and compare our results to those obtained in [2, 5].
Acknowledgement
The author is a JSPS International Research Fellow and his research is supported by JSPS KAKENHI Grant Number 23KF0114. He would like to thank Wenda Fang, Si-Qi Liu, Youjin Zhang and Chunhui Zhou for very helpful discussions.
2. Heisenberg vertex algebra and quantum Hamiltonian structure
2.1. Basics of vertex algebra
Let us review some basic facts about vertex algebras, and we refer to [8] for details. The paper [4] also serves as a good reference for the topic.
Recall that a vertex algebra is given by the data $\left(V,|0\rangle,T,Y(-,z)\right)$, where $V$ is a vector space over $\mathbb{C}$ called the space of states, $|0\rangle\in V$ is a special vector called the vacuum vector, $T\in\mathrm{End}(V)$ is a linear operator called the infinitesimal translation operator and $Y(-,z)\colon V\to\mathrm{End}(V)[[z^{\pm 1}]]$ is a linear morphism called the state-field correspondence. For each state $a\in V$, let us denote
$Y(a,z)=\sum_{n\in\mathbb{Z}}a_{(n)}z^{-n-1},\quad a_{(n)}\in\mathrm{End}(V).$
We call $a_{(n)}$ the $n$-th product by $a$. These data are required to satisfy the following axioms:
  • $|0\rangle_{(n)}=\delta_{n,-1}\mathrm{Id}_{V}$ and for any $a\in V$, the action of $a_{(n)}$ on the vacuum vector is given by
    $a_{(n)}|0\rangle=\begin{cases}0,&\text{for }n\geq 0,\\ \delta_{n,-1}a,&\text{for }n\leq-1.\end{cases}$
  • $T|0\rangle=0$ and for any $a\in V$, $[T,Y(a,z)]=\partial_{z}Y(a,z)$.
  • For any $a,b\in V$, there exists a positive integer $N\in\mathbb{Z}_{+}$ such that
    $(z-w)^{N}[Y(a,z),Y(b,w)]=0.$
A convenient tool for actual computation of vertex algebras is the so-called $\lambda$-bracket formalism ([3, 14]). For $a,b\in V$, we denote
$[a_{\lambda}b]=\sum_{n\geq 0}\frac{\lambda^{n}}{n!}a_{(n)}b.$
The $\lambda$-bracket satisfies the following properties, where many useful identities for vertex algebras are encoded:
  • Sesquilinearity:
    $[Ta_{\lambda}b]=-\lambda[a_{\lambda}b]$, and $[a_{\lambda}Tb]=(T+\lambda)[a_{\lambda}b]$.
  • Skewsymmetry:
    $[b_{\lambda}a]=-[a_{-\lambda-T}b]$.
  • Jacobi identity:
    $[a_{\lambda}[b_{\mu}c]]=[b_{\mu}[a_{\lambda}c]]+[[a_{\lambda}b]_{\lambda+\mu}c]$.
[…]
(2.1) $:ab:-:ba:=\int_{-T}^{0}d\lambda\,[a_{\lambda}b],$
(2.2) $::ab:c:-:a:bc::=:\left(\int_{0}^{T}d\lambda\,a\right)[b_{\lambda}c]:+:\left(\int_{0}^{T}d\lambda\,b\right)[a_{\lambda}c]:.$
To compute the right-hand sides of the above formulae, we must put $\lambda$ on the left under the sign of integral, and act the operators obtained from the definite integral to the right. For example, we have
[…]
When computing the $\lambda$-bracket of normal order product, we can use the following useful formula:
(2.3) $[a_{\lambda}:bc:]=:[a_{\lambda}b]c:+:b[a_{\lambda}c]:+\int_{0}^{\lambda}[[a_{\lambda}b]_{\mu}c]d\mu.$
Another standard fact about the normal product is that
(2.4) $Y(:ab:,z)=Y(a,z)_{+}Y(b,z)+Y(b,z)Y(a,z)_{-},$
where we set
$Y(a,z)_{+}=\sum_{n<0}a_{(n)}z^{-n-1},\quad Y(a,z)_{-}=\sum_{n\geq 0}a_{(n)}z^{-n-1}.$
In this paper, we will only consider the Heisenberg vertex algebra. Let $V$ be the Heisenberg vertex algebra of rank $N$ generated by states $b^{1},\dots,b^{N}$ satisfying the commutation relation
(2.5) $\left[b^{\alpha}_{(n)},b^{\beta}_{(m)}\right]=\hbar\,\eta^{\alpha\beta}\,n\,\delta_{m+n,0},\quad\alpha,\beta=1,\dots,N,$
where $(\eta^{\alpha\beta})$ is a constant non-degenerate symmetric matrix and $\hbar$ is just a formal parameter. The state space $V$ is a free $\mathbb{C}[\hbar]$-module spanned by states of the form
$b^{\alpha_{1}}_{(-k_{1}-1)}\dots b^{\alpha_{n}}_{(-k_{n}-1)}|0\rangle,\quad n,k_{i}\geq 0,\quad\alpha_{i}=1,\dots,N.$
The action of each $b^{\alpha}_{(n)}$ on state vectors is then uniquely determined by the relation (2.5). Note that in particular we have $b^{\alpha}_{(0)}=0$. Note that it follows from the vacuum axiom of $V$ that $b^{\alpha}=b^{\alpha}_{(-1)}|0\rangle$. For the Heisenberg vertex algebra, the state-field correspondence takes a very specific form, which can be described as follows.
Proposition 2.1 (Remark 2.2.6 of [8]).
For $n\geq 0$, $k_{1},\dots k_{n}\geq 0$ and $\alpha_{1},\dots,\alpha_{n}=1,\dots,N$, let us denote
$a=b^{\alpha_{1}}_{(-k_{1}-1)}\dots b^{\alpha_{n}}_{(-k_{n}-1)}|0\rangle\in V,$
then we have
$Y\left(a,z\right)=(-1)^{\sum k_{i}}\sum_{m_{1},\dots,m_{n}}\binom{m_{1}}{k_{1}}\dots\binom{m_{n}}{k_{n}}:b^{\alpha_{1}}_{(m_{1}-k_{1})}\dots b^{\alpha_{n}}_{(m_{n}-k_{n})}:z^{-n-\sum m_{i}},$
where the normal order on the right-hand side means that we always put annihilators, namely those operators $b^{\alpha}_{(m)}$ with $m\geq 0$, on the right.
Remark 2.1.
Note that the normal order defined above is well-defined due to the commutation relation (2.5). For general vertex algebras, we cannot define such normal order.
Proof.
Let us prove by induction on $n$. The case $n=0$ is trivial. Assume that we have proved the proposition for some $n\geq 0$, then let us consider the state
$a=b^{\alpha}_{(-k-1)}b^{\alpha_{1}}_{(-k_{1}-1)}\dots b^{\alpha_{n}}_{(-k_{n}-1)}|0\rangle\in V.$
First we notice that
$b^{\alpha}_{(-k-1)}=\frac{1}{k!}(T^{k}b^{\alpha})_{(-1)},$
hence we have
$a=\frac{1}{k!}:(T^{k}b^{\alpha})\left(b^{\alpha_{1}}_{(-k_{1}-1)}\dots b^{\alpha_{n}}_{(-k_{n}-1)}|0\rangle\right):.$
Then by using the fact that $Y(T^{k}b^{\alpha},z)=\partial_{z}^{k}Y(b^{\alpha},z)$ and the identity (2.4), we arrive at the following expression for $Y(a,z)$:
$\begin{split}(-1)^{k+\sum k_{i}}\sum_{\begin{subarray}{c}m_{1},\dots,m_{n}\\ m<0\end{subarray}}&\binom{m}{k}\binom{m_{1}}{k_{1}}\dots\binom{m_{n}}{k_{n}}\\ &\cdot b^{\alpha}_{(m-k)}:b^{\alpha_{1}}_{(m_{1}-k_{1})}\dots b^{\alpha_{n}}_{(m_{n}-k_{n})}:z^{-n-\sum m_{i}-m-1}\end{split}$
$\begin{split}+(-1)^{k+\sum k_{i}}\sum_{\begin{subarray}{c}m_{1},\dots,m_{n}\\ m\geq 0\end{subarray}}&\binom{m}{k}\binom{m_{1}}{k_{1}}\dots\binom{m_{n}}{k_{n}}\\ &\cdot:b^{\alpha_{1}}_{(m_{1}-k_{1})}\dots b^{\alpha_{n}}_{(m_{n}-k_{n})}:b^{\alpha}_{(m-k)}z^{-n-\sum m_{i}-m-1}.\end{split}$
In the first summation, note that $m-k<0$, hence by definition of the normal order, we have
$b^{\alpha}_{(m-k)}:b^{\alpha_{1}}_{(m_{1}-k_{1})}\dots b^{\alpha_{n}}_{(m_{n}-k_{n})}:\,=\,:b^{\alpha}_{(m-k)}b^{\alpha_{1}}_{(m_{1}-k_{1})}\dots b^{\alpha_{n}}_{(m_{n}-k_{n})}:,\quad m<0.$
In the second summation, note that for $10\leq m\leq k-$ 1, the binomial coefficient $\binom{m}{k}=0$, and for $m\geq k$ we have
$:b^{\alpha_{1}}_{(m_{1}-k_{1})}\dots b^{\alpha_{n}}_{(m_{n}-k_{n})}:b^{\alpha}_{(m-k)}=:b^{\alpha}_{(m-k)}b^{\alpha_{1}}_{(m_{1}-k_{1})}\dots b^{\alpha_{n}}_{(m_{n}-k_{n})}:.$
Therefore, we conclude that $Y(a,z)$ is indeed of the desired form and the proposition is proved. ∎
Corollary 2.2.
Fix a state $a=b^{\alpha_{1}}_{(-k_{1}-1)}\dots b^{\alpha_{n}}_{(-k_{n}-1)}|0\rangle$ for some $n\geq 0$ and $k_{i}\geq 0$, then we have
$a_{(m)}=\sum_{\begin{subarray}{c}j_{1}\leq\dots\leq j_{n}\\ \beta_{1},\dots,\beta_{n}\end{subarray}}C^{j_{1},\dots,j_{n}}_{\beta_{1},\dots,\beta_{n}}b^{\beta_{1}}_{(j_{1})}\dots b^{\beta_{n}}_{(j_{n})}\in\mathrm{End}(V),$
here $C^{j_{1},\dots,j_{n}}_{\beta_{1},\dots,\beta_{n}}$ are some constants (depending on the state $a$ and $m\in\mathbb{Z}$).
2.2. Quantum Hamiltonian structure
In this section, we construct a deformation quantization of Hamiltonian structure of hydrodynamic types using the Heisenberg vertex algebra. To this end, we will first recall the notion of a (classical) Hamiltonian structure, and one may refer to [16] for an introduction to the general theory of Hamiltonian structures.
Let $M$ be an $N$-dimensional smooth manifold, and let us denote by $\mathcal{A}$ the ring of differential polynomials of the infinite jet bundle $J^{\infty}M$. Locally, if we choose a coordinate system $(v^{1},\dots,v^{N})$ of $M$, then
$\mathcal{A}=\mathbb{C}\left[v^{\alpha,p}\colon\alpha=1,\dots,N,\,p\geq 0\right].$
Define a derivation $\partial_{x}$ of $\mathcal{A}$ by
$\partial_{x}=\sum_{p\geq 0}v^{\alpha,p+1}\frac{\partial}{\partial v^{\alpha,p}},$
here and henceforth we assume the summation over repeated lower and upper Greek indices. Then we denote by $\mathcal{F}=\mathcal{A}/\partial_{x}\mathcal{A}$ the space of local functionals, and for $f\in\mathcal{A}$, we denote by $\int f$ its class in $\mathcal{F}$. The ring $\mathcal{A}$ is graded by the differential degree defined by
$\deg_{\partial_{x}}v^{\alpha,p}=p,$
and we denote by $\mathcal{A}_{d}$ the space of homogeneous differential polynomials of differential degree $d$.
Remark 2.2.
In the literature, the ring of differential polynomials are defined by a certain completion of $\mathcal{A}$. For our purpose, this completion is not needed.
A Hamiltonian structure is a Lie algebra structure on $\mathcal{F}$ whose Lie bracket is of the following special form:
$\{F,G\}=\int\sum_{k\geq 0}\frac{\delta F}{\delta v^{\alpha}}P^{\alpha\beta}_{k}\partial_{x}^{k}\frac{\delta G}{\delta v^{\beta}},\quad P^{\alpha\beta}_{k}\in\mathcal{A},\quad\forall F,G\in\mathcal{F},$
here the variational derivatives are defined by
$\frac{\delta F}{\delta v^{\alpha}}=\sum_{p\geq 0}(-\partial_{x})^{p}\frac{\partial f}{\partial v^{\alpha,p}},\quad F=\int f.$
Let us consider the following special Hamiltonian structure of hydrodynamic type:
(2.6) $\{F,G\}=\int\frac{\delta F}{\delta v^{\alpha}}\eta^{\alpha\beta}\partial_{x}\frac{\delta G}{\delta v^{\beta}},$
where $(\eta^{\alpha\beta})$ is a non-degenerate symmetric constant matrix. It is proved in [9] that the bracket (2.6) serves as a universal local model for Hamiltonian structures with hydrodynamic leading terms. Such Hamiltonian structures are ubiquitous in the theory of integrable hierarchies. For example, all Dubrovin-Zhang hierarchies possess this kind of Hamiltonian structures.
Let us fix a Hamiltonian structure of the form (2.6) and let $V$ be the Heisenberg vertex algebra of rank $N$ generated by states $b^{1},\dots,b^{N}$ satisfying
$\left[b^{\alpha}_{(n)},b^{\beta}_{(m)}\right]=\hbar\,\eta^{\alpha\beta}\,n\,\delta_{m+n,0},\quad\alpha,\beta=1,\dots,N.$
The key observation is that we can identify the state space and the space of differential polynomials via the $\hbar$-linear morphism
(2.7) $\varphi:\mathcal{A}[\hbar]\to V\colon v^{\alpha_{1},k_{1}}\dots v^{\alpha_{n},k_{n}}\mapsto k_{1}!\dots k_{n}!b^{\alpha_{1}}_{(-k_{1}-1)}\dots b^{\alpha_{n}}_{(-k_{n}-1)}|0\rangle.$
Lemma 2.3.
$\varphi$ is an isomorphism between vector spaces, and for any $f\in\mathcal{A}[\hbar]$, we have
(2.8) $\varphi\left(\partial_{x}f\right)=T\varphi\left(f\right),$
and for any $p\geq 0$,
(2.9) $\varphi\left(v^{\alpha,p}f\right)=p!\,b^{\alpha}_{(-p-1)}\varphi(f),\quad\hbar\,\eta^{\alpha\beta}\varphi\left(\frac{\partial f}{\partial v^{\beta,p}}\right)=\frac{1}{(p+1)!}b^{\alpha}_{(p+1)}\varphi(f).$
Proof.
These facts are straightforward to verify, and we omit the details. ∎
It follows from (2.8) that $\varphi$ induces an isomorphism between quotient spaces $\mathcal{F}[\hbar]$ and $V/TV$, and we will still denote this isomorphism by $\varphi$. For general vertex algebras, it is well-known that $V/TV$ admits a natural Lie algebra structure given by the 00-th product.
Lemma 2.4 (Theorem 4.1.2 of [8]).
Let $\pi\colon V\to V/TV$ be the projection map. Then the bracket
(2.10) $[A,B]=\pi\left(\pi^{-1}(B)_{(0)}\pi^{-1}(A)\right),\quad A,B\in V/TV$
is well-defined, and defines a Lie algebra structure on $V/TV$.
Via $\varphi$, the above bracket on $V/TV$ can be identified to give a bracket on $\mathcal{F}[\hbar]$. Our first goal is to compute this bracket in terms of differential polynomials, and hence prove that this bracket gives a deformation quantization of the Hamiltonian structure (2.6). For this purpose, let us compute $\varphi(f)_{(n)}$ for $f\in\mathcal{A}$ and $n\geq 0$.
Recall that $V$ is a free $\mathbb{C}[\hbar]$-module, hence we can decompose
$V=\bigoplus_{n\geq 0}V_{n},\quad V_{n}=\hbar^{n}V_{0},$
where $V_{0}$ is the $\mathbb{C}$-vector space spanned by states of the form $b^{\alpha_{1}}_{(-k_{1}-1)}\dots b^{\alpha_{n}}_{(-k_{n}-1)}|0\rangle$. Then it is obvious that for $a\in V_{0}$, we have $(\hbar^{m}a)_{(n)}=\hbar^{m}\cdot a_{(n)}$, and we can decompose
$a_{(n)}=\sum_{i\geq 0}a_{(n)}^{[i]},\quad a_{(n)}^{[i]}\colon V_{0}\to V_{i}.$
By using Corollary 2.2, if we denote
$a_{(n)}=\sum_{j_{1}\leq\dots\leq j_{n}}C^{j_{1},\dots,j_{n}}_{\beta_{1},\dots,\beta_{n}}b^{\beta_{1}}_{(j_{1})}\dots b^{\beta_{n}}_{(j_{n})},$
then it follows from the commutation relation (2.5) that
$a_{(n)}^{[i]}=\sum_{(j_{1},\dots,j_{n})\in J^{i}}C^{j_{1},\dots,j_{n}}_{\beta_{1},\dots,\beta_{n}}b^{\beta_{1}}_{(j_{1})}\dots b^{\beta_{n}}_{(j_{n})},$
where we use $J^{i}$ to denote the index set
$\{(j_{1},\dots,j_{n})\in\mathbb{Z}^{n}\colon j_{1}\leq\dots\leq j_{n-i}\leq 0\leq j_{n-i+1}\leq\dots\leq j_{n}\}.$
Lemma 2.5.
Fix a state $a=b^{\alpha_{1}}_{(-k_{1}-1)}\dots b^{\alpha_{n}}_{(-k_{n}-1)}|0\rangle\in V_{0}$ for some $n\geq 0$ and $k_{i}\geq 0$. Then we have
$a_{(-1)}^{[0]}=b^{\alpha_{1}}_{(-k_{1}-1)}\dots b^{\alpha_{n}}_{(-k_{n}-1)}.$
Proof.
Using Proposition 2.1, we have
$a_{(-1)}=(-1)^{\sum k_{i}}\sum_{m_{1}+\dots+m_{n}=-n}\binom{m_{1}}{k_{1}}\dots\binom{m_{n}}{k_{n}}:b^{\alpha_{1}}_{(m_{1}-k_{1})}\dots b^{\alpha_{n}}_{(m_{n}-k_{n})}:,$
and by the above observation, we arrive at
$a_{(-1)}^{[0]}=(-1)^{\sum k_{i}}\sum_{\begin{subarray}{c}m_{1}+\dots+m_{n}=-n,\\ m_{i}\leq k_{i}\end{subarray}}\binom{m_{1}}{k_{1}}\dots\binom{m_{n}}{k_{n}}:b^{\alpha_{1}}_{(m_{1}-k_{1})}\dots b^{\alpha_{n}}_{(m_{n}-k_{n})}:.$
In order to obtain a non-vanishing summand, we must have $m_{i}\leq-1$, and therefore the only non-vanishing summand is given by the term with indices $m_{1}=\dots=m_{n}=-1$, namely,
$a_{(-1)}^{[0]}=b^{\alpha_{1}}_{(-k_{1}-1)}\dots b^{\alpha_{n}}_{(-k_{n}-1)}.$
The lemma is proved. ∎
Remark 2.3.
This lemma also follows directly from the fact that $a_{(-1)}|0\rangle=a$ for any $a\in V$.
We have the following immediate consequence.
Corollary 2.6.
For $f,g\in\mathcal{A}$, we have
$\varphi(f)_{(-1)}^{[0]}\varphi(g)=\varphi(fg).$
Now we are ready to prove the main result of this section.
Theorem 2.7.
Let $f,g\in\mathcal{A}$ be two differential polynomials. Then for $n\geq 0$, we have
$\varphi^{-1}\left(\varphi(f)_{(n)}\varphi(g)\right)=\sum_{m\geq 1}\frac{\hbar^{m}}{m!}$ $\sum_{\begin{subarray}{c}r_{1},\dots,r_{m}\geq 0\\ s_{1},\dots,s_{m}\geq 0\\ 2m-1+\sum r_{i}+s_{i}\geq n\end{subarray}}\frac{(-1)^{\sum r_{i}}\prod(r_{i}+s_{i}+1)!\eta^{\alpha_{i}\beta_{i}}}{(2m-1-n+\sum r_{i}+\sum s_{i})!}$
$\cdot\frac{\partial^{m}g}{v^{\beta_{1},s_{1}}\dots v^{\beta_{m},s_{m}}}\partial_{x}^{2m-1-n+\sum r_{i}+\sum s_{i}}\left(\frac{\partial^{m}f}{\partial v^{\alpha_{1},r_{1}}\dots\partial v^{\alpha_{m},r_{m}}}\right).$
Proof.
By a similar argument as in the proof of Lemma 2.5, we see that $\varphi(f)_{(n)}^{[0]}=0$ for $n\geq 0$. Therefore, we only need to consider $\varphi(f)_{(n)}^{[m]}$ for $m\geq 1$. Without loss of generality, we may write
$\varphi(f)_{(n)}^{[m]}=\sum_{(j_{1},\dots,j_{\ell})\in J^{m}}C^{j_{1},\dots,j_{\ell}}_{\beta_{1},\dots,\beta_{\ell}}b^{\beta_{1}}_{(j_{1})}\dots b^{\beta_{\ell}}_{(j_{\ell})},\quad C^{j_{1},\dots,j_{\ell}}_{\beta_{1},\dots,\beta_{\ell}}\in\mathbb{C}.$
It is more convenient to rewrite this operator into the form
$\varphi(f)_{(n)}^{[m]}=\sum_{k_{1},\dots,k_{m}>0}A^{k_{1},\dots,k_{m}}_{\alpha_{1},\dots,\alpha_{m}}b^{\alpha_{1}}_{(k_{1})}\dots b^{\alpha_{m}}_{(k_{m})},\quad A^{k_{1},\dots,k_{m}}_{\alpha_{1},\dots,\alpha_{m}}\in\mathrm{End}(V),$
here $A^{k_{1},\dots,k_{m}}_{\alpha_{1},\dots,\alpha_{m}}$ are operators given by linear combinations of creators of the form
$b^{\beta_{1}}_{(-j_{1}-1)}\dots b^{\beta_{\ell-m}}_{(-j_{\ell-m}-1)},\quad j_{1},\dots,j_{\ell-m}\geq 0,$
and in particular, we may and do fix each operator $A^{k_{1},\dots,k_{m}}_{\alpha_{1},\dots,\alpha_{m}}$ uniquely by requiring that it is symmetric with respect to indices $(\alpha_{1},k_{1}),\dots,(\alpha_{m},k_{m})$. Therefore, by using (2.9), we arrive at
$\varphi(f)_{(n)}^{[m]}\varphi(g)=\hbar^{m}\sum_{k_{1},\dots,k_{m}>0}$ $k_{1}!\dots k_{m}!\,\eta^{\alpha_{1}\gamma_{1}}\dots\eta^{\alpha_{m}\gamma_{m}}$
$\cdot A^{k_{1},\dots,k_{m}}_{\alpha_{1},\dots,\alpha_{m}}\varphi\left(\frac{\partial^{m}g}{\partial v^{\gamma_{1},k_{1}-1}\dots\partial v^{\gamma_{m},k_{m}-1}}\right).$
In particular, it follows that
$\varphi(f)_{(n)}^{[m]}\varphi(v^{\beta_{1},s_{1}-1}\dots v^{\beta_{m},s_{m}-1})=$ $\hbar^{m}\sum_{k_{1},\dots,k_{m}>0}k_{1}!\dots k_{m}!\,\eta^{\alpha_{1}\gamma_{1}}\dots\eta^{\alpha_{m}\gamma_{m}}$
$\cdot A^{k_{1},\dots,k_{m}}_{\alpha_{1},\dots,\alpha_{m}}\sum_{\sigma\in S_{m}}\prod_{i=1}^{m}\delta^{\beta_{\sigma(i)}}_{\gamma_{i}}\delta_{k_{i},s_{\sigma(i)}}|0\rangle$
=\,= $\hbar^{m}m!\,s_{1}!\dots s_{m}!\,\eta^{\alpha_{1}\beta_{1}}\dots\eta^{\alpha_{m}\beta_{m}}A^{s_{1},\dots,s_{m}}_{\alpha_{1},\dots,\alpha_{m}}|0\rangle.$
Note that we have used the fact that $A^{k_{1},\dots,k_{m}}_{\alpha_{1},\dots,\alpha_{m}}$ is symmetric with respect to indices. Hence, by using Lemma 2.5, we have
$\varphi(f)_{(n)}^{[m]}\varphi(g)=\frac{\hbar^{m}}{m!}\sum_{s_{1},\dots,s_{m}\geq 0}$ $\left(\varphi(f)_{(n)}^{[m]}\varphi(v^{\beta_{1},s_{1}}\dots v^{\beta_{m},s_{m}})\right)_{(-1)}^{[0]}$
$\cdot\varphi\left(\frac{\partial^{m}g}{\partial v^{\beta_{1},s_{1}}\dots\partial v^{\beta_{m},s_{m}}}\right).$
It follows from the skew-symmetry property of the $\lambda$-bracket that for general $A,B\in V$,
$A_{(n)}B=\sum_{k\geq 0}\frac{(-1)^{k+n+1}}{k!}T^{k}\left(B_{(k+n)}A\right),$
hence we only need to compute
$\varphi(v^{\beta_{1},s_{1}}\dots v^{\beta_{m},s_{m}})_{(k+n)}^{[m]}\varphi(f),\quad k\geq 0.$
By using the definition (2.7) of $\varphi$ and Proposition 2.1, we have
$\varphi($ $v^{\beta_{1},s_{1}}\dots v^{\beta_{m},s_{m}})_{(k+n)}^{[m]}=(-1)^{\sum s_{i}}s_{1}!\dots s_{m}!$
$\sum_{\begin{subarray}{c}n_{1},\dots,n_{m}\geq 0\\ \sum n_{i}+\sum s_{i}+2m=k+n+1\end{subarray}}\binom{n_{1}+s_{1}+1}{s_{1}}\dots\binom{n_{m}+s_{m}+1}{s_{m}}b^{\beta_{1}}_{(n_{1}+1)}\dots b^{\beta_{m}}_{(n_{m}+1)}.$
Then after a straightforward computation, we arrive at
$\varphi(f)_{(n)}^{[m]}\varphi(v^{\beta_{1},s_{1}}\dots v^{\beta_{m},s_{m}})$
$=\sum_{\begin{subarray}{c}r_{1},\dots,r_{m}\geq 0\\ \sum r_{i}+\sum s_{i}+2m\geq n+1\end{subarray}}\frac{(-1)^{\sum r_{i}}\prod_{i=1}^{m}(r_{i}+s_{i}+1)!\eta^{\alpha_{i}\beta_{i}}}{\left(\sum r_{i}+\sum s_{i}+2m-n-1\right)!}$
$\cdot T^{\sum r_{i}+\sum s_{i}+2m-n-1}\left(\varphi\left(\frac{\partial^{m}f}{\partial v^{\alpha_{1},r_{1}}\dots\partial v^{\alpha_{m},r_{m}}}\right)\right).$
Finally, we arrive at the desired formula by using the identity (2.8) and Corollary 2.6. The theorem is proved. ∎
Recall that we have a Lie bracket on $V/TV$ defined by the formula (2.10). By identifying $V/TV$ with $\mathcal{F}[\hbar]$ via $\varphi$, it follows from the above theorem that we have a Lie bracket on $\mathcal{F}[\hbar]$ defined by
$[F,G]=\int\sum_{m\geq 1}\frac{\hbar^{m}}{m!}$ $\sum_{\begin{subarray}{c}r_{1},\dots,r_{m}\geq 0\\ s_{1},\dots,s_{m}\geq 0\end{subarray}}\frac{(-1)^{\sum r_{i}}\prod(r_{i}+s_{i}+1)!\eta^{\alpha_{i}\beta_{i}}}{(2m-1+\sum r_{i}+\sum s_{i})!}$
$\cdot\frac{\partial^{m}f}{v^{\beta_{1},s_{1}}\dots v^{\beta_{m},s_{m}}}\partial_{x}^{2m-1+\sum r_{i}+\sum s_{i}}\left(\frac{\partial^{m}g}{\partial v^{\alpha_{1},r_{1}}\dots\partial v^{\alpha_{m},r_{m}}}\right),$
where $f,g\in\mathcal{A}[\hbar]$ are arbitrary lifts of $F$ and $G$, respectively. This Lie bracket defines a deformation quantization of the Hamiltonian structure (2.6) in the sense that
(2.11) $\lim_{\hbar\to 0}\frac{1}{\hbar}[F,G]=\{F,G\},\quad F,G\in\mathcal{F}.$
In particular, we have proved Theorem 1.1 and Proposition 1.4 as special cases of Theorem 2.7 for $N=1$ and $\eta=1$. As an application of the above formula, let us prove the following interesting fact which will be used later.
Proposition 2.8.
Let $a\in V$ be a state with $a_{(0)}=0$. Then there exist polynomials $B_{\alpha}(\hbar),C(\hbar)\in\mathbb{C}[\hbar]$ such that
$a=B_{\alpha}(\hbar)b^{\alpha}+C(\hbar)|0\rangle+TV.$
Proof.
Let us denote by $f=\varphi^{-1}(a)\in\mathcal{A}[\hbar]$, and we decompose
$f=\sum_{n\geq 0}\hbar^{n}f_{n},\quad f_{n}\in\mathcal{A}.$
The fact that $a_{(0)}=0$ implies that, for any $g\in\mathcal{A}$, $\varphi(f)_{(0)}\varphi(g)=0$. Then by taking the $\hbar$-coefficient of $\varphi(f)_{(0)}\varphi(g)$, we arrive at
(2.12) $\sum_{r,s\geq 0}\eta^{\alpha\beta}(-1)^{r}\frac{\partial g}{\partial v^{\beta,s}}\partial_{x}^{r+s+1}\frac{\partial f_{0}}{\partial v^{\alpha,r}}=0.$
In particular, we have
$\{G,F_{0}\}=\int\frac{\delta G}{\delta v^{\alpha}}\eta^{\alpha\beta}\partial_{x}\frac{\delta F_{0}}{\delta v^{\beta}}=0,\quad F_{0}=\int f_{0},\quad G=\int g.$
Then it follows from Lemma 2.1.7 of [17] that there exist constants $A$ and $B^{\alpha}$ such that
$\eta^{\alpha\beta}\frac{\delta f_{0}}{\delta v^{\beta}}=Av^{\alpha}+B^{\alpha},$
which implies that there exists another constant $C$ such that
$f_{0}=\frac{1}{2}\eta_{\alpha\beta}Av^{\alpha}v^{\beta}+\eta_{\alpha\beta}B^{\alpha}v^{\beta}+C+\partial_{x}\mathcal{A},$
here we denote by $\eta_{\alpha\beta}$ the inverse of $\eta^{\alpha\beta}$. Due to (2.12), we see that $A=0$, hence we conclude that
$f_{0}=\eta_{\alpha\beta}B^{\alpha}v^{\beta}+C+\partial_{x}\mathcal{A},$
which implies that actually $\varphi(f_{0})_{(0)}=0$. By passing to $(f-f_{0})/\hbar$, the same argument shows that $f_{1}$ is of the same form as $f_{0}$. In this way, we prove that actually each $f_{n}$ has the same form of $f_{0}$, namely, there exist polynomials $B_{\alpha}(\hbar),C(\hbar)\in\mathbb{C}[\hbar]$ such that
$f=B_{\alpha}(\hbar)v^{\alpha}+C(\hbar)+\partial_{x}\mathcal{A}[\hbar].$
The proposition is proved. ∎
3. Weyl quantization of dispersionless KdV hierarchy
In this section, let us consider the quantization of the dispersionless KdV hierarchy. To state the problem in a precise way, recall that the classical dispersionless KdV hierarchy is a family of compatible PDEs
$\frac{\partial v}{\partial t_{n}}=\frac{v^{n}}{n!}\frac{\partial v}{\partial x},\quad n\geq 0$
for the unknown function $v$. This system admits a Hamiltonian structure given by
$\{F,G\}=\int\frac{\delta F}{\delta v}\partial_{x}\frac{\delta G}{\delta v},\quad F,G\in\mathcal{F},$
and each flow $\frac{\partial}{\partial t_{n}}$ is the Hamiltonian vector filed with Hamiltonian $H_{n+1}^{c}$, where
$H_{n}^{c}=\int\frac{v^{n+1}}{(n+1)!},\quad n\geq 0.$
It follows that these Hamiltonians are in involution:
$\{H_{n}^{c},H_{m}^{c}\}=0,\quad n,m\geq 0.$
To consider the quantization problem, we replace the Hamiltonian structure by the deformed bracket. Then the classical Hamiltonians do not commute anymore. For example, we have
$[H_{2}^{c},H_{3}^{c}]=\int\frac{\hbar}{4}v^{4}v_{x}+\frac{\hbar^{2}}{12}(3vv_{x}v_{xx}+v^{2}v_{xxx})+\frac{\hbar^{3}}{720}v^{(5)}\neq 0.$
Therefore, the quantization problem is to find a quantum correction $H_{n}$ of $H_{n}^{c}$ such that
$[H_{n},H_{m}]=0,\quad\lim_{\hbar\to 0}H_{n}=H_{n}^{c}.$
3.1. Construction of quantum Hamiltonians
From now on, let us denote by $V$ the vertex algebra of rank $=1$. It is generated by the field
$Y(b,z)=\sum_{n\in\mathbb{Z}}b_{(n)}z^{-n-1},\quad\left[b_{(n)},b_{(m)}\right]=\hbar\,n\,\delta_{m+n}.$
Our idea to construct the quantum Hamiltonians is as follows. The ring $\mathcal{A}$ of differential polynomials is a commutative algebra, and we can view the normal order product on $V$ as a deformation of this algebra structure, in the sense that
$\lim_{\hbar\to 0}:\varphi(f)\varphi(g):=\varphi(fg),\quad f,g\in\mathcal{A},$
here recall that $:ab:=a_{(-1)}b$ for $a,b\in V$. The normal product is neither commutative nor associative, and we can use the idea of Weyl quantization in quantum mechanics to quantize the Hamiltonians $H_{n}^{c}$. Roughly speaking, in Weyl quantization, a quantum Hamiltonian is obtained from a classical Hamiltonian by averaging all the possible ways to compose all the quantum operators appeared in Hamiltonian. For example, let $(x,p)$ be a pair of conjugated variables, then we have the following Weyl quantization:
$(xp)^=\frac{1}{2}(\hat{x}\hat{p}+\hat{p}\hat{x}),\quad(xp^{2})^=\frac{1}{3}(\hat{x}\hat{p}\hat{p}+\hat{p}\hat{x}\hat{p}+\hat{p}\hat{p}\hat{x}).$
In our case, we quantize the Hamiltonians $H_{n}^{c}$ by averaging all the possible ways to associate the state $b=\varphi(v)$ using the normal order product. For example, we should define
$h_{2}=\frac{1}{12}\varphi^{-1}(:b:bb::+::bb:b:),\quad H_{2}=\int h_{2}.$
We call this procedure a non-associative Weyl quantization.
To give a precise definition of the quantum Hamiltonians, let us define the set $BT(n)$ of labeled full binary rooted tress with $n$ leaves for $n\geq 1$ recursively. Each element in $BT(n)$ will be a full binary tree whose nodes, except for the root node, are labeled either by $L$ (for left) or $R$ (for right). In addition, $n$ leaves of a tree in $BT(n)$ are labeled by $1,\dots,n$. $BT(1)$ is just the set of a single root node and $BT(2)$ is the set of the unique full binary tree with two leaves labeled by $L,1$ and $R,2$.
RootL,1R,2
Figure 1. The unique tree in $BT(2)$.
For each tree $T\in BT(n)$ and for a choice of its leaf with label $i\in\{1,\dots,n\}$, let us define a new tree $T_{i}$ as follows. Firstly we make this leaf with label $i$ an internal node by adding to it 2 leaves, and hence obtain a full binary tree with $n+1$ leaves. Then the label ‘$i$’ of this node is removed. For leaves originally labeled by $11,\dots,i-$ 1 their labels remain unchanged, and for leaves originally labeled by $i+1,\dots,n$ their labels are added by one. Finally, the newly added leaves are labeled as $L,i$ and $R,i+1$. We then define
$BT(n+1):=\coprod_{T\in BT(n)}\{T_{i}\mid i=1,\dots,n\}.$
From the construction, it is easy to see that $|BT(n)|=(n-1)!$. As an example, elements in $BT(3)$ is presented in Fig. 2.
RootLL,1R,2R,3
RootL,1RL,2R,3
Figure 2. Trees in $BT(3)$. If we denote by $T$ the unique tree in $BT(2)$, then the first tree is $T_{1}$ and the second one is $T_{2}$.
Let us proceed to associate a state to each tree in $BT(n)$. For each tree $T\in BT(n)$, we first put at every leaf the state $b\in V$, and then we recursively put at every other nodes the state obtained by taking the normal order product of states associated with its left child and right child. More explicitly, given a node which is not a leaf, assume that we have put $f\in V$ at its left child and $g\in V$ at its right child, then this node is associated with the state $:fg:\in V$. We denote by $A_{T}\in V$ the state associated with the root of $T$. For example, for $T_{1},T_{2}\in BT(3)$ as presented in Fig. 2, we have
$A_{T_{1}}=::bb:b:,\quad A_{T_{2}}=:b:bb::.$
[…]
Figure 3. Evaluation of each node for trees in $BT(3)$.
Let us denote by $Z_{n}$ the state obtained by averaging all the states $A_{T}$ for $T\in BT(n)$, namely, we define
$Z_{n}=\frac{1}{(n-1)!}\sum_{T\in BT(n)}A_{T}.$
The following Lemma is a straightforward consequence of our definition.
Lemma 3.1.
For $n\geq 2$, we have
$Z_{n}=\frac{1}{n-1}\sum_{i=1}^{n-1}:Z_{i}Z_{n-i}:.$
By using Corollary 2.6, it is easy to see that
$\lim_{\hbar\to 0}\varphi^{-1}(Z_{n})=v^{n},$
hence let us define the quantum Hamiltonian densities
$h_{n}=\frac{1}{(n+1)!}\varphi^{-1}(Z_{n+1}),\quad n\geq 0.$
Alternatively, we can equivalently define $h_{n}$ by the recursion relation
(3.1) $h_{0}=\varphi^{-1}(b),\quad h_{n}=\frac{1}{n}\sum_{k=0}^{n-1}\frac{1}{\binom{n+1}{k+1}}\varphi^{-1}\left(:\varphi(h_{k})\varphi(h_{n-k-1}):\right),\quad n\geq 1.$
Example 3.1.
It is obvious that $h_{0}=v$ and $h_{1}=\frac{v^{2}}{2}$. Let us compute $h_{2}$. First we see
$:b:bb::=b_{(-1)}b_{(-1)}b_{(-1)}|0\rangle,$
and by using the quasi-associativity (2.2) of normal order product, we have
$::bb:b:=:b:bb::+:\left(\int_{0}^{T}d\lambda\,b\right)[b_{\lambda}b]:=:b:bb::+\hbar T^{2}b,$
here we use the fact that
$[b_{\lambda}b]=\sum_{n\geq 0}\frac{\lambda^{n}}{n!}b_{(n)}b_{(-1)}|0\rangle=\hbar\lambda|0\rangle.$
Therefore, it follows that
$Z_{3}=\frac{1}{2}\left(::bb:b:+:b:bb::\right)=b^{3}+\frac{\hbar}{2}T^{2}b,\quad b^{3}=:b:bb::,$
and hence
$h_{2}=\frac{1}{6}v^{3}+\frac{\hbar}{12}v_{xx}.$
Example 3.2.
Let us compute $h_{3}$. We begin by enumerate elements in $BT(4)$. Denote by $R$ and $S$ the first and the second tree presented in Fig 2, then by definition we arrive at
$BT(4)=\{R_{1},R_{2},R_{3},S_{1},S_{2},S_{3}\}.$
These trees are presented in the following diagrams.
RootLLL,1R,2R,3R,4
RootLL,1RL,2R,3R,4
RootLL,1R,2RL,3R,4
Figure 4. Trees $R_{1}$, $R_{2}$ and $R_{3}$.
RootLL,1R,2RL,3R,4
RootL,1RLL,2R,3R,4
RootL,1RL,2RL,3R,4
Figure 5. Trees $S_{1}$, $S_{2}$ and $S_{3}$.
Let us compute the contribution from each graph. We start from
$A_{R_{3}}=A_{S_{1}}=:b^{2}b^{2}:,\quad b^{2}=\,:bb:.$
By using formula (2.2), it follows that
$:b^{2}b^{2}:-b^{4}=2\left(\int_{0}^{T}d\lambda\,b\right)[b_{\lambda}b^{2}]=2\hbar:T^{2}bb:,\quad b^{4}=\,:b:b:bb:::,$
here we use (2.3) to obtain
$[b_{\lambda}b^{2}]=:[b_{\lambda}b]b:+:b[b_{\lambda}b]:+\int_{0}^{\lambda}[[b_{\lambda}b]_{\mu}b]d\mu=2\hbar\lambda b.$
From the above computation we see that
$A_{R_{3}}=A_{S_{1}}=b^{4}+2\hbar:T^{2}bb:.$
Next it is easy to obtain that
$A_{S_{2}}=:b:b^{2}b::=:b\left(b^{3}+\hbar T^{2}b\right):\,=b^{4}+\hbar:bT^{2}b:.$
Note that it follows from the quasi-commutativity (2.1) that
$:bT^{2}b:\,=\,:T^{2}bb:.$
To compute $A_{R_{2}}$, we may use the following quasi-commutativity formula:
$b^{4}-:b^{3}b:=\int_{-T}^{0}d\lambda[b_{\lambda}b^{3}]=\int_{-T}^{0}d\lambda(3\hbar\lambda b^{2})=-3\hbar(:TbTb:+:bT^{2}b:).$
Hence, we arrive at
$A_{R_{2}}=:b^{3}b:=b^{4}+3\hbar(:TbTb:+:bT^{2}b:).$
Finally, we have $A_{S_{3}}=b^{4}$ and
$A_{R_{1}}=::b^{2}b:b:=:(b^{3}+\hbar T^{2}b)b:=b^{4}+3\hbar:TbTb:+4\hbar:bT^{2}b:.$
By averaging the contribution from each tree, we have
$Z_{4}=b^{4}+\hbar:TbTb:+2\hbar:bT^{2}b:,$
which gives
$h_{3}=\frac{v^{4}}{24}+\frac{\hbar}{24}\left(v_{x}^{2}+2vv_{xx}\right).$
Example 3.3.
By a similar but more complicated computation, we compute $h_{4}$ by averaging contributions from 24 trees and $h_{5}$ by averaging 120 trees, and the result is
$h_{4}$ $=\frac{v^{5}}{120}+\frac{\hbar}{24}\left(vv_{x}^{2}+v^{2}v_{xx}\right)+\frac{\hbar^{2}}{360}v^{(4)},$
$h_{5}$ $=\frac{v^{6}}{720}+\frac{\hbar}{72}\left(\frac{3}{2}v^{2}v_{x}^{2}+v^{3}v_{xx}\right)+\frac{\hbar^{2}}{360}\left(vv^{(4)}+2v_{x}v_{xxx}+\frac{3}{2}v_{xx}^{2}\right).$
Lemma 3.2.
We have the following homogeneous properties for $h_{n}$:
$\left(\sum_{k\geq 0}(k+1)v^{(k)}\frac{\partial}{\partial v^{(k)}}\right)h_{n}$ $=(n+1)h_{n},$
$\left(-2\hbar\frac{\partial}{\partial\hbar}+\sum_{k\geq 0}kv^{(k)}\frac{\partial}{\partial v^{(k)}}\right)h_{n}$ $=0.$
Proof.
Note that by definition, each $Z_{n}$ is of conformal degree $n$, where we define each $b_{(n)}$ to be of conformal degree $-n$. This conformal degree of $V$ induces a gradation on $\mathcal{A}[\hbar]$ via $\varphi$, with $v^{(k)}$ to be of degree $k+1$, hence we prove the first identity.
For the second identity, we are to consider the differential degree of $h_{n}$. We define the differential degree of $\mathcal{A}[\hbar]$ to be
$\deg_{\partial_{x}}v^{(n)}=n,\quad\deg_{\partial_{x}}\hbar=-2.$
This gradation induces a gradation on $V$, where $\deg_{\partial_{x}}b_{(n)}=-n-1$. For $f\in\mathcal{A}[\hbar]$ of degree $d$, it is easy to see, using Proposition 2.1 that
$\deg_{\partial_{x}}\varphi(f)_{(n)}=d-n-1.$
Then we can prove by induction, using the recursion (3.1), that each $h_{n}$ is of differential degree 00. The lemma is proved. ∎
3.2. Another expression for quantum Hamiltonians
As we have already seen, it is not easy to compute each $h_{n}$ using the definition. In what follows, we give another expression of $h_{n}$, which makes it easy to compute $h_{n}$, and more importantly, enables us to prove various properties of quantum Hamiltonians. From now on, for simplicity, we will set $\hbar=1$, and the dependence of $\hbar$ may be recovered from homogeneous conditions.
We start by defining states $p^{(n)}\in V$ for $n\geq 0$:
$p^{(0)}=b,\quad p^{(n+1)}=\frac{1}{n+2}\left(Tp^{(n)}+:bp^{(n)}:\right),\quad n\geq 0.$
Denote by $g_{n}=\varphi^{-1}(p^{(n)})\in\mathcal{A}$, then by using relations (2.8) and (2.9), it is easy to see that $g_{n}$ satisfies the recursion relation
$g_{0}=v,\quad g_{n+1}=\frac{1}{n+2}\left(\partial_{x}g_{n}+vg_{n}\right),\quad n\geq 0.$
This recursion relation makes it easy to compute each $g_{n}$, and in particular, it follows that $g_{n}$ are given by Schur polynomials. Recall that for a partition $\lambda=(\lambda_{1},\dots,\lambda_{\ell(\lambda)})$, we associate it with a differential polynomial $S_{\lambda}\in\mathcal{A}$, called the Schur polynomial associated with the partition $\lambda$, by
$S_{\lambda}=\det{\left({\bf{e}}_{\lambda_{i}-i+j}\right)_{1\leq i,j\leq\ell(\lambda)}},$
where ${\bf{e}}_{n}\in\mathcal{A}$ is given by
$\sum_{n\in\mathbb{Z}}{\bf{e}}_{n}z^{n}=\exp\left(\sum_{k\geq 0}\frac{v^{(k)}}{(k+1)!}z^{k+1}\right).$
Lemma 3.3.
We have $g_{n}=S_{(n+1)}$ for $n\geq 0$.
Proof.
Let us denote by
$G(z)=\sum_{n\geq 0}g_{n}z^{n}$
the generating function of $g_{n}$, then we see that this function is uniquely determined by the equation
$z\frac{\partial G}{\partial z}=z(\partial_{x}G+vG)-G+v,\quad G(0)=v.$
Then it suffices to show that
$G(z)=\frac{W}{z}-\frac{1}{z},\quad W=\exp\left(\sum_{k\geq 0}\frac{v^{(k)}}{(k+1)!}z^{k+1}\right).$
Indeed, it follows that
$\partial_{x}\left(\frac{W}{z}-\frac{1}{z}\right)=\frac{1}{z}\sum_{k\geq 0}v^{(k+1)}\frac{\partial W}{\partial v^{(k)}}=\frac{W}{z}\left(\sum_{k\geq 0}\frac{v^{(k+1)}}{(k+1)!}z^{k+1}\right),$
$z\frac{\partial}{\partial z}\left(\frac{W}{z}-\frac{1}{z}\right)=\frac{\partial W}{\partial z}-\frac{W}{z}+\frac{1}{z}=vW+\partial_{x}W-\frac{W}{z}+\frac{1}{z}.$
Then the lemma is proved by a straightforward computation. ∎
Remark 3.1.
We will give another proof of the above lemma in the next subsection based on the definition of Hall-Littlewood polynomials.
The differential polynomials $g_{n}$ are closely related to $h_{n}$. In fact, we have the following key observation which makes it possible to prove various properties of $h_{n}$.
Theorem 3.4.
For $n\geq 0$, we have
$h_{n}=\sum_{k=0}^{n}\frac{(-1)^{n-k}}{(n-k+1)!}\partial_{x}^{n-k}g_{k}.$
We will prove this theorem in the next subsection.
3.3. Properties of quantum Hamiltonians
Let us prove various properties of quantum Hamiltonian densities $h_{n}$ of dispersionless KdV hierarchy. Our main technique is based on the vertex operator representation of W-type operators and their relations to Hall-Littlewood polynomials, which is studied in [12]. These operators have a deep relationship with and find wide application in the theory of (classical) integrable hierarchy such as KP theory and its reduction (see. e.g., [18] and references therein), and we are now to apply these techniques to study the quantum integrable hierarchy.
We start by making necessary preparations. Denote by
$J(z)=\sum_{n\geq 0}\frac{v^{(n)}}{n!}z^{n}+\sum_{n\geq 0}(n+1)!\frac{\partial}{\partial v^{(n)}}z^{-n-2}\in\mathrm{End}(\mathcal{A})[[z^{\pm 1}]],$
then we see that $J(z)$ is precisely
[…]
Let us then define
$P^{(k)}(z)=\frac{1}{(k+1)!}:(\partial_{z}+J(z))^{k}J(z):=\sum_{m\in\mathbb{Z}}P^{(k)}_{m}z^{-k-m-1},$
here the normal order product means to put the annihilators, i.e. the operators $\frac{\partial}{\partial v^{(n)}}$ on the right. Then it follows that
[…]
in particular, it follows from the relation $\varphi(g_{k})_{(-1)}|0\rangle=\varphi(g_{k})$ that
$g_{k}=P^{(k)}_{-k-1}(1).$
Denote by
$\phi(z)=\sum_{n\geq 1}\frac{v^{(n-1)}}{n!}z^{n}-\sum_{n\geq 1}(n-1)!\frac{\partial}{\partial v^{(n-1)}}z^{-n},$
then we have ([18])
(3.2) $P^{(k)}(z)=\frac{1}{(k+1)!}:e^{-\phi(z)}\partial_{z}^{k+1}e^{\phi(z)}:.$
We define $B_{n},B^{*}_{n}\in\mathrm{End}(\mathcal{A})$ by
$B(z)=:e^{\phi(z)}:=\sum_{n\in\mathbb{Z}}B_{n}z^{n},\quad B^{*}(z)=:e^{-\phi(z)}:=\sum_{n\in\mathbb{Z}}B_{n}^{*}z^{-n},$
these operators satisfy the relation ([12])
(3.3) $B_{m}B_{n}+B_{n-1}B_{m+1}$ $=0,$
(3.4) $B_{m}^{*}B_{n}^{*}+B_{n+1}^{*}B_{m-1}^{*}$ $=0,$
(3.5) $B_{m}B_{n}^{*}+B_{n-1}^{*}B_{m-1}$ $=\delta_{m,n}.$
We can represent each operator $P^{(k)}_{m}$ via $B_{n}$ and $B_{n}^{*}$ by using (3.2).
Lemma 3.5 (Theorem 3.3 in [18]).
The following relation holds true:
$P^{(k)}_{m}=-\sum_{n\in\mathbb{Z}}\binom{n}{k}B^{*}_{m+n}B_{n}.$
Operators $B_{n}$ can be used to define the Schur polynomials ([19]). Given a partition $\lambda=(\lambda_{1},\dots,\lambda_{\ell(\lambda)})$, it follows that
[…]
furthermore, the following relation also holds true (identity (72) of [18]):
$B^{*}_{-n}(1)=(-1)^{n}S_{(\underbrace{1,\dots,1}_{n})}.$
It is much more convenient to work with Hall-Littlewood polynomials, namely for any $(\lambda_{1},\dots,\lambda_{\ell})\in\mathbb{Z}^{\ell}$, we define
[…]
Note that for $\lambda_{1}\geq\lambda_{2}\dots\geq\lambda_{\ell}>0$, the corresponding Hall-Littlewood polynomial coincides with the Schur polynomial.
Using the above ingredients, for $n\geq 0$, we arrive at
(3.6) $P^{(\ell)}_{-n-1}(1)$ $=-\sum_{a\in\mathbb{Z}}\binom{a}{\ell}B^{*}_{a-n-1}B_{a}(1)$
$=\sum_{a=0}^{n}\binom{a}{\ell}B_{a+1}B_{a-n}^{*}(1)$
$=\sum_{a=0}^{n}(-1)^{a-n}\binom{a}{\ell}B_{a+1}S_{(\underbrace{1,\dots,1}_{n-a})}$
$=\sum_{a=0}^{n}(-1)^{a-n}\binom{a}{\ell}S_{(a+1,\underbrace{1,\dots,1}_{n-a})},$
here for the second equality we use the relation (3.5) and the fact that $B^{*}_{k}(1)=0$ for $k>0$. In particular, by setting $\ell=n$, it follows that
$P^{(n)}_{-n-1}(1)=S_{(n+1)},$
which gives another proof of Lemma 3.3.
Now we are ready to study properties of quantum Hamiltonians. Our strategy is as follows: first we define $\tilde{h}_{n}\in\mathcal{A}$ by
(3.7) $\tilde{h}_{n}=\sum_{k=0}^{n}\frac{(-1)^{n-k}}{(n-k+1)!}\partial_{x}^{n-k}g_{k},$
then we prove Proposition 1.2, Theorem 1.3, Theorem  1.5 and Proposition 1.6 with $h_{n}$ replaced by $\tilde{h}_{n}$. Finally, we prove that $\tilde{h}_{n}$ actually satisfies the recursion relation (3.1), hence $\tilde{h}_{n}=h_{n}$, and this proves Theorem 3.4.
3.3.1. Proposition 1.6 for $\tilde{h}_{n}$
We start by expressing $\tilde{h}_{n}$ via Schur polynomials. It follows from the definition (3.7) that
[…]
then we arrive at the relation
$\tilde{h}_{n}=\sum_{k=0}^{n}\frac{(-1)^{k}}{k+1}P^{(n-k)}_{-n-1}(1).$
Applying the identity (3.6), it turns out that
(3.9) $\tilde{h}_{n}$ $=\sum_{k=0}^{n}\frac{(-1)^{k}}{k+1}\sum_{a=n-k}^{n}(-1)^{a-n}\binom{a}{n-k}S_{(a+1,\underbrace{1,\dots,1}_{n-a})}$
$=\frac{1}{n+1}\sum_{k=0}^{n}\frac{1}{\binom{n}{k}}S_{(n-k+1,\underbrace{1,\dots,1}_{k})},$
here we use the following binomial identity proved in Corollary 2.2 of [22]:
$\sum_{k=0}^{n}\frac{(-1)^{k}}{m+k+1}\binom{n}{k}=\frac{1}{(n+m+1)\binom{n+m}{m}},\quad n,m\in\mathbb{Z}_{\geq 0}.$
In this way, we prove Proposition 1.6 for $\tilde{h}_{n}$.
3.3.2. Proposition 1.2 for $\tilde{h}_{n}$
Let us proceed to prove that differential polynomials $\tilde{h}_{n}$ satisfy the recursion relations (1.5). Denote by
[…]
then it follows from Proposition 1.4 that
$\mathcal{D}=\sum_{s,t\geq 0}\frac{(s+t+1)!}{s!t!}v^{(s)}v^{(t)}\frac{\partial}{\partial v^{(s+t)}}+\frac{(s+1)!(t+1)!}{(s+t+2)!}v^{(s+t+2)}\frac{\partial^{2}}{\partial v^{(s)}\partial v^{(t)}}.$
Equivalently, using (3.8), this operator can also be written as
$\mathcal{D}=P^{(1)}_{-1}+2P^{(2)}_{-1},$
and we need to prove that
$\mathcal{D}(\tilde{h}_{n})=(n+1)(n+2)\tilde{h}_{n+1},\quad n\geq 0.$
The actions of $P^{(k)}_{m}$ on Hall-Littlewood polynomials are computed in [18]. Precisely, fix $(\lambda_{1},\dots,\lambda_{\ell})\in\mathbb{Z}^{\ell}$, then
(3.10) $P^{(k)}_{m}(S_{\lambda})=$ $\sum_{i=1}^{\ell}\binom{\lambda_{i}-m-i}{k}S_{(\lambda_{1},\dots,\lambda_{i}-m,\dots,\lambda_{\ell})}+\delta_{m,0}\binom{-\ell}{k+1}S_{\lambda}$
$+\delta_{m<0}\sum_{i=1}^{-m}(-1)^{m-i}\binom{i-\ell-1}{k}S_{(\lambda_{1},\dots,\lambda_{\ell},i,\underbrace{1,\dots,1}_{-m-i})},$
where we set $\delta_{m<0}=1$ for $m<0$ and $\delta_{m<0}=0$ for $m\geq 0$. Let us use this to compute $\mathcal{D}(\tilde{h}_{n})$. Due to (3.9), let us first consider the action
$P^{(1)}_{-1}S_{(n-k+1,\underbrace{1,\dots,1}_{k})}$
for $0\leq k\leq n$. Using the relation (3.3), we have $B_{1}B_{2}=0$, hence
$S_{(n-k+1,1,\dots,1,2,1,\dots,1)}=0.$
Then it follows that
$P^{(1)}_{-1}S_{(n-k+1,\underbrace{1,\dots,1}_{k})}=(n-k+1)S_{(n-k+2,\underbrace{1,\dots,1}_{k})}-(k+1)S_{(n-k+1,\underbrace{1,\dots,1}_{k+1})},$
and similarly we have
$2P^{(2)}_{-1}S_{(n-k+1,\underbrace{1,\dots,1}_{k})}=(n-k+1)(n-k)S_{(n-k+2,\underbrace{1,\dots,1}_{k})}+(k+1)(k+2)S_{(n-k+1,\underbrace{1,\dots,1}_{k+1})}.$
Then it follows that
$\mathcal{D}(\tilde{h}_{n})=$ $\,\frac{1}{n+1}\sum_{k=0}^{n}\frac{1}{\binom{n}{k}}\mathcal{D}(S_{(n-k+1,\underbrace{1,\dots,1}_{k})})$
$=$ $\,\frac{1}{n+1}\sum_{k=0}^{n}\frac{(n-k+1)^{2}}{\binom{n}{k}}S_{(n-k+2,\underbrace{1,\dots,1}_{k})}$
$+\frac{1}{n+1}\sum_{k=0}^{n}\frac{(k+1)^{2}}{\binom{n}{k}}S_{(n-k+1,\underbrace{1,\dots,1}_{k+1})}$
$=$ $\,(n+1)\sum_{k=0}^{n+1}\frac{1}{\binom{n+1}{k}}S_{(n-k+2,\underbrace{1,\dots,1}_{k})}$
$=$ $\,(n+1)(n+2)\tilde{h}_{n+1}.$
This proves Proposition 1.2 for $\tilde{h}_{n}$.
3.3.3. Proposition 1.5 for $\tilde{h}_{n}$
Let us consider the eigenvalue problems for $\tilde{h}_{n}$. By using (3.10), we see that all the Schur polynomials $S_{\lambda}$ serve as common eigenvectors of $P^{(k)}_{0}$. It follows from (3.8) that
[…]
Then we prove Proposition 1.5 for $\tilde{h}_{n}$ by applying (3.10). Furthermore, we recover the $\hbar$-dependence by noticing that $\varphi(\tilde{h}_{n})_{(n)}$ is of differential degree $-n-1$ and $S_{\lambda}$ is of differential degree zero.
3.3.4. Theorem 1.3 for $\tilde{h}_{n}$
Next let us prove that the functionals $\tilde{H}_{n}=\int\tilde{h}_{n}$ are in involution with respect to the deformed quantum Hamiltonian structure, which is equivalent to prove
$\varphi(\tilde{h}_{n})_{(0)}\varphi(\tilde{h}_{m})\in TV,\quad m,n\geq 0.$
Proposition 3.6.
For $m,n\geq 0$, we have
$\left[\varphi(\tilde{h}_{n})_{(0)},\varphi(\tilde{h}_{m})_{(0)}\right]=0.$
Proof.
By using the relation (3.8), it suffices to show that
$\left[P^{(n)}_{-n},P^{(m)}_{-m}\right]=0,\quad m,n\geq 0.$
It follows from Lemma 3.5 that
$\left[P^{(n)}_{-n},P^{(m)}_{-m}\right]=\,$ $\sum_{a,b\in\mathbb{Z}}\binom{a}{m}\binom{b}{n}\left[B^{*}_{a-n}B_{a},B^{*}_{b-m}B_{b}\right]$
=\,= $\sum_{a,b\in\mathbb{Z}}\binom{a}{m}\binom{b}{n}B^{*}_{a-n}B_{a}B^{*}_{b-m}B_{b}-B^{*}_{b-m}B_{b}B^{*}_{a-n}B_{a}$
=\,= $\sum_{a\in\mathbb{Z}}\binom{a}{n}\binom{a+m}{m}B^{*}_{a-n}B_{a+m}-\sum_{a\in\mathbb{Z}}\binom{a}{n}\binom{a-n}{m}B^{*}_{a-n-m}B_{a}$
=\,= $0,$
here for the third equality, we use the relations (3.3) –(3.5). The proposition is proved. ∎
By applying the well-known identity
(3.11) $[A_{(n)},B_{(m)}]=\sum_{i\geq 0}\binom{n}{i}(A_{(i)}B)_{m+n-i},\quad A,B\in V,\quad n,m\in\mathbb{Z},$
we conclude that
(3.12) $\left(\varphi(\tilde{h}_{n})_{(0)}\varphi(\tilde{h}_{m})\right)_{(0)}=0,$
hence it follows from Proposition 2.8 that there exist constants $B_{n,m}$ and $C_{n,m}$ such that
$\varphi(\tilde{h}_{n})_{(0)}\varphi(\tilde{h}_{m})=B_{n,m}b+C_{n,m}|0\rangle+TV.$
We notice that each $\varphi(\tilde{h}_{n})$ is of conformal degree $n+1$, hence $\varphi(\tilde{h}_{n})_{(0)}\varphi(\tilde{h}_{m})$ is of conformal degree $n+m+1$, which implies that
$\varphi(\tilde{h}_{n})_{(0)}\varphi(\tilde{h}_{m})\in TV.$
This proves Theorem 1.3 for $\tilde{h}_{n}$.
3.3.5. Proof of Theorem (3.4)
Finally, we are to prove $\tilde{h}_{n}=h_{n}$. Due to $\tilde{h}_{0}=h_{0}$, it suffices to verify the recursion relation
$\varphi(\tilde{h}_{n+1})=\frac{1}{n+1}\sum_{k=0}^{n}\frac{1}{\binom{n+2}{k+1}}\varphi(\tilde{h}_{k})_{(-1)}\varphi(\tilde{h}_{n-k}),\quad n\geq 0.$
Let us prove the validity of the above relation by induction on $n$. The case $n=0$ can be directly verified, and we assume that we have proved this for $n=0,1,\dots,N-1$ for some $N\geq 1$. Now let us consider the case $n=N$. Firstly, it follows from the recursion
$\mathcal{D}(\tilde{h}_{N})=(N+1)(N+2)\tilde{h}_{N+1}$
that
$\varphi(\tilde{h}_{N+1})=\frac{2}{(N+1)(N+2)}\varphi(\tilde{h}_{2})_{(1)}\varphi(\tilde{h}_{N}).$
Therefore, by using the induction hypothesis, we arrive at
$\varphi(\tilde{h}_{N+1})$ $=\frac{2}{(N+1)(N+2)}\varphi(\tilde{h}_{2})_{(1)}\varphi(\tilde{h}_{N})$
$=\frac{2}{N(N+1)(N+2)}\sum_{k=0}^{N-1}\frac{1}{\binom{N+1}{k+1}}\varphi(\tilde{h}_{2})_{(1)}\varphi(\tilde{h}_{k})_{(-1)}\varphi(\tilde{h}_{N-1-k}).$
By combining (3.11) and (3.12), we arrive at
$\left[\varphi(\tilde{h}_{2})_{(1)},\varphi(\tilde{h}_{k})_{(-1)}\right]=\left(\varphi(\tilde{h}_{2})_{(1)}\varphi(\tilde{h}_{k})\right)_{(-1)}.$
Finally, we can compute straightforwardly as
$\varphi(\tilde{h}_{N+1})=\,$ $\frac{2}{N(N+1)(N+2)}\sum_{k=0}^{N-1}\frac{1}{\binom{N+1}{k+1}}\varphi(\tilde{h}_{2})_{(1)}\varphi(\tilde{h}_{k})_{(-1)}\varphi(\tilde{h}_{N-1-k})$
=\,= $\frac{2}{N(N+1)(N+2)}\sum_{k=0}^{N-1}\frac{1}{\binom{N+1}{k+1}}\varphi(\tilde{h}_{k})_{(-1)}\varphi(\tilde{h}_{2})_{(1)}\varphi(\tilde{h}_{N-1-k})$
$+\frac{2}{N(N+1)(N+2)}\sum_{k=0}^{N-1}\frac{1}{\binom{N+1}{k+1}}\left(\varphi(\tilde{h}_{2})_{(1)}\varphi(\tilde{h}_{k})\right)_{(-1)}\varphi(\tilde{h}_{N-1-k})$
$=$ $\,\frac{1}{N(N+1)(N+2)}\sum_{k=0}^{N-1}\frac{(N-k)(N-k+1)}{\binom{N+1}{k+1}}\varphi(\tilde{h}_{k})_{(-1)}\varphi(\tilde{h}_{N-k})$
$+\frac{1}{N(N+1)(N+2)}\sum_{k=0}^{N-1}\frac{(k+1)(k+2)}{\binom{N+1}{k+1}}\varphi(\tilde{h}_{k+1})_{(-1)}\varphi(\tilde{h}_{N-1-k})$
$=$ $\frac{1}{N+1}\sum_{k=0}^{N}\frac{1}{\binom{N+2}{k+1}}\varphi(\tilde{h}_{k})_{(-1)}\varphi(\tilde{h}_{N-k}),$
and we conclude that $\tilde{h}_{n}=h_{n}$.
4. Conclusion
In this paper, we study the relationship between vertex algebras and quantum integrable hierarchies. This idea is quite natural, indeed, Barakat, De Sole, Kac and their collaborators developed a theory to describe Hamiltonian integrable hierarchies via Poisson vertex algebras (see. e.g., [13]), and these algebras, in some cases, can be viewed as a classical limit of vertex algebras. In this sense, the vertex algebras should provide suitable tools to study quantum integrable hierarchies, and we investigate this possibility in this paper via the example of quantum dispersionless KdV hierarchy.
Let us compare the results of the present paper to those obtained in the papers [2, 5]. In the first place, a different deformation quantization is constructed in [2], which is given by the commutator of the standard Moyal product corresponding to the Hamiltonian structure. In terms of differential polynomials, the deformed Hamiltonian structure reads
$[F,G]=\int\sum_{m\geq 1}\frac{\hbar^{m}}{m!}$ $\sum_{\begin{subarray}{c}r_{1},\dots,r_{m}\geq 0\\ s_{1},\dots,s_{m}\geq 0\end{subarray}}\frac{\partial^{m}f}{v^{\beta_{1},s_{1}}\dots v^{\beta_{m},s_{m}}}\prod_{i=1}^{m}(-1)^{r_{i}}\eta^{\alpha_{i}\beta_{i}}$
×$\times\sum_{j=1}^{2m-1+\sum r_{i}+\sum s_{i}}C_{j}^{r_{1}+s_{1}+1,\dots,r_{m}+s_{m}+1}\partial_{x}^{j}\left(\frac{\partial^{m}g}{\partial v^{\alpha_{1},r_{1}}\dots\partial v^{\alpha_{m},r_{m}}}\right),$
where $C_{j}^{a_{1},\dots,a_{m}}$ are certain constants, and in particular we have
$C^{a_{1},\dots,a_{m}}_{m-1+\sum a_{i}}=\frac{a_{1}!\dots a_{m}!}{(m-1+\sum a_{i})!}.$
Therefore, the above bracket can be viewed as a further deformation of the bracket constructed in the present paper.
Another difference is that, in the construction of [2, 5], what matters is the quantum Hamiltonians, instead of quantum Hamiltonian densities. Indeed, the eigenvalue problem considered in [5] only depends on the deformed Hamiltonians as local functionals, and does not depend on the choice of densities. However, in our construction, the densities $h_{n}$ are important since we are considering operators
[…]
and we cannot add arbitrary terms in $\partial_{x}\mathcal{A}$ to $h_{n}$.
Despite these differences, our results also share similarity to the known results. Therefore, it makes sense to conjecture that our formulation is equivalent to the formulation presented in [2, 5] in some sense. For example, can we actually recover the results in [5] using results of the present paper? Is the above bracket constructed in [2] related to the bracket presented in this paper by a certain transformation? These problems are still under investigation.